# Graded Betti numbers of the Jacobian algebra and total Tjurina numbers of plane curves

### v2

@@ 2. Graded Betti numbers and partitions
with $c_{1}\leq...\leq c_{m-2}$ and $d_{1}\leq...\leq d_{m}$. It follows from [17, Lemma 1.1] that
@@
Using [17, Formula (13)], one obtains the relation
@@
see for instance [6, inequality (2.6)]. In this note we prove the following inequality
(2.8) $E=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}\geq 0,$
which is a new necessary condition on these Betti numbers. In other words, (2.8) is not a consequence of the inequalities (2.5), (2.6) and (2.7), see for details Remark 5.3 below.
@@ 3. Tjurina numbers for curves via graded Betti numbers of their Jacobian algebra
(i) Theorem 3.1 is a generalization of [2, Proposition 1.11] which treats the case $t=2$ and of [13, Theorem 3.1] which treats the case $t=3$.
@@
see for instance [7]. Since in this case $t(C)=0$, we see that the equality in Theorem 3.1 holds in this case as well with the obvious conventions
@@ 4. Examples for curves of type $t(C)=4$
The following is a direct consequence of Theorem 3.1. Consider all the possible partitions of $t$, namely:
@@
We recall first the following result due to du Plessis and Wall, see [14, Theorem 3.2] as well as [16] for an alternative approach.
@@
We recall that a maximal Tjurina curve $C$ is a curve such that $2d_{1}\geq d$ and $\tau(C)=\tau(d,d_{1})^{\prime}_{max}$, where see [10]. As an example, the curve in Example 4.2 (viii) is a maximal Tjurina curve with $d=5$ and $d_{1}=4$. Theorem 3.1 yields a precise way to measure the difference between $\tau(C)$ and $\tau(d,d_{1})_{max}$ (resp. $\tau(d,d_{1})^{\prime}_{max}$), and in this way gives, in particular, a proof of the claims involving these invariants in Theorem 5.1.
@@
Moreover, all the terms in the right hand sides in (1) and (2) are positive.
@@
$E=\tau(C)-\tau(d,d_{1})_{min}=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}.$
@@
Indeed, $d_{1}-t=d-1-d_{2}\geq 0$ and $d-1-d_{3}\geq 0$ by (2.6).
On the other hand, for $m\geq 4$, the inequality $E\geq 0$ coming from Theorem 5.1 is a new necessary condition to be satisfied by the exponents $(d_{1},d_{2},\ldots,d_{m})$ and $(\epsilon_{1},\dots,\epsilon_{n})$, the corresponding ordered partition of $t$.
To show that this inequality is not a consequence of the inequalities (2.5), (2.6) and (2.7), one may consider the data $d_{1}=3$, $d_{2}=d_{3}=8$, $d_{4}=15$, $\epsilon_{1}=\epsilon_{2}=1$ which verify the inequalities (2.5), (2.6) and (2.7), but not the above inequality. Indeed, for these data one gets E=1<0E=-1<0, which shows that there is no reduced curve $C$ with these Betti numbers.
Recall that an $3$-syzygy curve $C$ is called nearly free (resp. minimal plus-one generated) if $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$ (resp. $d_{1}+d_{2}=d$ and $d_{2}+1=d_{3}$, see [8, 11]. Using Corollary 5.2 (1) we can give new proofs for the following known results. Indeed, the claim (1) below occurs in [5, Theorem 1.3], and the claim (2) below occurs in [11, Theorem 1.5].
@@
Using Corollary 5.2 (2) we can give a completely new proof of [10, Theorem 3.1], that characterizes the maximal Tjurina curves in terms of their resolutions.
@@
Let $C:f=0$ be a reduced plane curve of degree $d$ and type $t(C)=t\geq 1$ with exponents $(d_{1},d_{2},\ldots,d_{m})$ and $\pi_{C}=(\epsilon_{1},\dots,\epsilon_{n})$ the corresponding ordered partition of $t$, with n=m2n=m-2. Then $C$ is a nearly maximal Tjurina curve if and only if one of the following two situations occurs.
@@
from [12, Remark 4.3]. Then $C$ is a 4-syzygy curve with exponents $(6,6,6,7)$ and degree $d=11$. It follows that
@@
(ii) Using the final part of Remark 5.7 we have the following. The curves $C$ and $C^{\prime}$ in [12, Example 4.5] satisfy the conditions in Corollary 5.6 (2) for $d=7$, $d_{1}=4$ and $m=3$. The curves $C$ in [12, Example 4.6] satisfies the conditions in Corollary 5.6 (2) for $d=9$, $d_{1}=5$ and $m=3$. The curves $C$ in [12, Example 4.7] satisfies the conditions in Corollary 5.6 (2) for $d=11$, $d_{1}=6$ and $m=3$. To get examples of curves satisfying the conditions in Corollary 5.6 (2) with $m\geq 4$, one may consider the curves $C$, $^{\prime}, C′′′C^{\prime\prime\prime}$ and $C_{6}$ in [12, Example 5.1], the curves $C_{7}$ and $C^{\prime}$ in [12, Example 5.2], the curves $C$ and $C_{8}$ in [12, Example 5.3] and the curves $C^{\prime}$ and $C^{\prime\prime\prime}$ in [12, Example 5.4].

### v3

@@ 2. Graded Betti numbers and partitions
with $c_{1}\leq...\leq c_{m-2}$ and $d_{1}\leq...\leq d_{m}$. It follows from [16, Lemma 1.1] that
@@
Using [16, Formula (13)], one obtains the relation
@@
see for instance [6, inequality (2.6)]. Let us consider the following expression involving the graded Betti numbers
(2.8) $E=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2},$
where $t=t(C)$ as in (2.4). In this note we prove the following inequality
(2.9) $E\geq 0$
with equality if and only if $m=3$ and $d_{2}=d_{3}=d-1$, which is a new necessary condition to be satisfied by the graded Betti numbers of a reduced plane curve. In other words, (2.9) is not a consequence of the inequalities (2.5), (2.6) and (2.7), see for details Remark 5.3 below.
@@ 3. Tjurina numbers for curves via graded Betti numbers of their Jacobian algebra
(i) Theorem 3.1 is a generalization of [2, Proposition 1.11] which treats the case $t=2$ and of [12, Theorem 3.1] which treats the case $t=3$.
@@
see for instance Theorem 5.1 below. Since in this case $t(C)=0$, we see that the equality in Theorem 3.1 holds in this case as well with the obvious conventions
@@ 4. Examples for curves of type $t(C)=4$
The following is a direct consequence of Theorem 3.1. Consider all the possible partitions of $t=4$, namely:
@@
We recall first the following result due to du Plessis and Wall, see [13, Theorem 3.2] as well as [15] for an alternative approach.
@@
We recall that a maximal Tjurina curve $C$ is a curve such that $2d_{1}\geq d$ and $\tau(C)=\tau(d,d_{1})^{\prime}_{max}$, see [9]. As an example, the curve in Example 4.2 (viii) is a maximal Tjurina curve with $d=5$ and $d_{1}=4$. Theorem 3.1 yields a precise way to measure the difference between $\tau(C)$ and $\tau(d,d_{1})_{max}$ (resp. $\tau(d,d_{1})^{\prime}_{max}$), and in this way gives, in particular, a proof of the claims involving these invariants in Theorem 5.1.
@@
Moreover, all the terms in the sums in right hand sides in (1) and (2) are positive.
@@
$E:=\tau(C)-\tau(d,d_{1})_{min}=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}.$
@@
Indeed, $d_{1}-t=d-1-d_{2}\geq 0$ and $d-1-d_{3}\geq 0$ by (2.6). In particular, the equality $E=0$ is equivalent to
$d_{2}=d_{3}=d-1$
when $m=3$. On the other hand, for $m\geq 4$, the inequality $E\geq 0$ coming from Theorem 5.1 is a new necessary condition to be satisfied by the exponents $(d_{1},d_{2},\ldots,d_{m})$ and $(\epsilon_{1},\dots,\epsilon_{n})$, the corresponding ordered partition of $t$. In fact using [8, Theorem 3.5 (1)] we get the following more precise inequality for the graded Betti numbers of any reduced curve $C$
(5.1) $E>0$
unless $m=3$ and $d_{2}=d_{3}=d-1$ when $E=0$. To show that this inequality is not a consequence of the inequalities (2.5), (2.6) and (2.7), one may consider the data $d_{1}=3$, $d_{2}=d_{3}=8$, $d_{4}=15$, $\epsilon_{1}=\epsilon_{2}=1$ which verify the inequalities (2.5), (2.6) and (2.7), but not the above inequality. Indeed, for these data one gets E=1<0E=-1<0, which shows that there is no reduced curve $C$ with these Betti numbers.
Recall that an $3$-syzygy curve $C$ is called nearly free (resp. minimal plus-one generated) if $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$ (resp. $d_{1}+d_{2}=d$ and $d_{2}+1=d_{3}$, see [7, 10]. Using Corollary 5.2 (1) we can give new proofs for the following known results. Indeed, the claim (1) below occurs in [5, Theorem 1.3], and the claim (2) below occurs in [10, Theorem 1.5].
@@
Using Corollary 5.2 (2) we can give a completely new proof of [9, Theorem 3.1], that characterizes the maximal Tjurina curves in terms of their resolutions.
@@
Let $C:f=0$ be a reduced plane curve of degree $d$ and type $t(C)=t\geq 1$ with exponents $(d_{1},d_{2},\ldots,d_{m})$ and $\pi_{C}=(\epsilon_{1},\dots,\epsilon_{n})$ the corresponding ordered partition of $t$, where n=m2n=m-2. Then $C$ is a nearly maximal Tjurina curve if and only if one of the following two situations occurs.
@@
from [11, Remark 4.3]. Then $C$ is a 4-syzygy curve with exponents $(6,6,6,7)$ and degree $d=11$. It follows that
@@
(ii) Using the final part of Remark 5.7 we have the following. The curves $C$ and $C^{\prime}$ in [11, Example 4.5] satisfy the conditions in Corollary 5.6 (2) for $d=7$, $d_{1}=4$ and $m=3$. The curves $C$ in [11, Example 4.6] satisfies the conditions in Corollary 5.6 (2) for $d=9$, $d_{1}=5$ and $m=3$. The curves $C$ in [11, Example 4.7] satisfies the conditions in Corollary 5.6 (2) for $d=11$, $d_{1}=6$ and $m=3$. To get examples of curves satisfying the conditions in Corollary 5.6 (2) with $m\geq 4$, one may consider the curves $C$, $^{\prime}, C′′′C^{\prime\prime\prime}$ and $C_{6}$ in [11, Example 5.1], the curves $C_{7}$ and $C^{\prime}$ in [11, Example 5.2], the curves $C$ and $C_{8}$ in [11, Example 5.3] and the curves $C^{\prime}$ and $C^{\prime\prime\prime}$ in [11, Example 5.4].
Conflict of Interests
We declare that there is no conflict of interest regarding the publication of this paper.
Data Availability Statement
We do not analyze or generate any datasets, because this work proceeds within a theoretical and mathematical approach.